We need to check the pedestals for L10B2 west hybrids 1 and 2. The software from last year didn't show them swapped but Dave noticed they were in the hardware so switched the cables. Need to check that everything is OK this year.
From March 6th to March 7th Seq Hi 0 Seq Lo 2 Thr Hi 7 Thr Lo 3
From March 7th to end of run Seq Hi 0 Seq Lo 2 Thr Hi 7 Thr Lo 2
Currently, we also write 1 pre and 4 post timebins.
New list Feb 11 from Run 404200
Feb 10th no longer write out anode 2 black
Feb 11th from run 4042007 no longer write out anode 239 black
Jan 24th T0 ~ 8 Dave moved timing so T0 now moves
Feb 5th To moved on 1 RHIC tick so T0 now at 11 ( Bill Christie's note: As of yesterday, the data taken with the modified trigger tier one files, with the bits in the trigger for most of the trigger detectors delayed by one STAR clock tick to line up with the BEMC trigger bits, looked good. Good in this sense means that the trigger data (ADCs, TACs, etc.), as well as the TPC and FTPC data, looked the same as the data taken with the tier one files we've been running with to date.)
West 5 back in as of Feb 4.